Grade 12th passIIT JEE Entrance Exam

A cane filled with water is revolved in a vertical circle of radius 4 m and the water just does not fall down the time period of revolution will be

Harsh
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
at the top most position
three forces will act on the water
1 weight in the downward direction
2 normal reaction by bucket in the downward direction
3 centrifugal force in the upward direction
so mg + N =mv2/r
for water just does not fall   N ≥ 0
so mg = mv2/r
v =√gr
so time period T = 2∏r/v
=2*3.14*4 /√4g
=3.14*4 /√9.8
= 4.01